# Efficient parallel scheduling for sparse triangular solvers

### v1

@@ 3. Barrier list schedulers
1 ${\rm superstep}\leftarrow 1$
2 ${\rm endStep}\leftarrow\textnormal{{false}}$
3 ${\rm finishTimes}\leftarrow\{0\}$
4 ${\rm free}[p]\leftarrow\textnormal{{true}}$, for all $p\in P$
5 ${\rm prio}[v],{\rm prio}[p,v]\leftarrow$ initialize priority, for all $v\in V$ and $p\in P$
6 ${\rm ready}_{\rm pool},\,{\rm ready}_{\rm all}\leftarrow\text{{sources of }}G$
7 ${\rm ready}_{p}\leftarrow\emptyset$, for all $p\in P$
8 while $\exists$ unassigned vertex do
9       if ${\rm endStep}$ and ${\rm finishTimes}=\emptyset$ then
10             ${\rm ready}_{p}\leftarrow\emptyset$, for all $p\in P$
11             ${\rm ready}_{\rm all}\leftarrow{\rm ready}_{\rm pool}$
12             ${\rm superstep}\leftarrow{\rm superstep}+1$
13             ${\rm endStep}\leftarrow\textnormal{{false}}$
14             ${\rm finishTimes}\leftarrow\{0\}$
@@
16      $t\leftarrow$ earliest time from ${\rm finishTimes}$
17       for all vertices $v$ that finish at $t$ do
18             ${\rm free}[\pi(v)]\leftarrow\textnormal{{true}}$
19             for all children $u$ of $v$ do
20                   if all parents of $u$ are finished then
21                         ${\rm ready}_{\rm pool}.{\rm insert}(u)$
22                         if $\forall\,(w,u)\in E:(\pi(w)=\pi(v)\text{ or }\sigma(w)<{\rm superstep})$ then
23                               ${\rm ready}_{\pi(v)}.{\rm insert}(u)$
24
25
@@ 18             ${\rm free}[\pi(v)]\leftarrow\textnormal{{true}}$
28      if $\neg{\rm endStep}$ or fill-up possible then
29             while $\exists p$ with ${\rm free}[p]=\textnormal{{true}}$ and ${\rm ready}_{p}\cup{\rm ready}_{\rm all}\neq\emptyset$ do
30                   if $\neg{\rm endStep}$ then
31                         $p,\,v\leftarrow$ choose using priority
32
33                   else
34                         $p,\,v\leftarrow$ choose using fill-up strategy
35
36                  delete $v$ from ${\rm ready}_{\rm pool}$, ${\rm ready}_{p}$, and ${\rm ready}_{\rm all}$
37                   $\pi(v)\leftarrow p$
38                   $\sigma(v)\leftarrow{\rm superstep}$
39                   ${\rm finishTimes}.{\rm insert}(t+w(v))$
40                   ${\rm free}[p]\leftarrow\textnormal{{false}}$
                   update priority in $v$’s neighborhood
                    // only in Locking
41
42
43      if $\alpha$ fraction of cores idle and ${\rm ready}_{\rm pool}$ large enough  then
44             ${\rm endStep}\leftarrow\textnormal{{true}}$
45
46      ${\rm finishTimes}.{\rm remove}(t)$
47
48return $(\pi,\sigma)$
Algorithm 3.1 Skeleton of our scheduling algorithms
  • In the $p$-ivotal path scheduler, we attempt to prioritize vertices that have long outgoing paths and/or a high number of successors in the (vertex-weighted) DAG $G=(V,E,\omega)$. For this, we use a static scoring function. The score of each vertex $v\in V$ is computed recursively using the parameter $p\geq 1$:
    ${\rm prio}[v]=\omega(v)+\left(\sum_{u\in{\rm Children}(v)}{\rm prio}[u]^{p}\right)^{\frac{1}{p}}.$ (3.1)
    Note that choosing $p\to\infty$ is just the longest path to a sink vertex excluding communication, which is sometimes referred to as bottom (vertex) distance. In our implementation, we have simply chosen $p=2$. We remark that one needs to be wary of numerical instability (overflow) in the implementation of this priority.
  • In the Locking scheduler, we aim to specifically avoid vertices that would lock out some other vertex from the current superstep due to a communication requirement. In particular, the assignment of vertex $v$ to a core $p$ gets a penalty value of $\ell$ if there are $\ell$ yet-uncomputed children of $v$ that already have parents computed on (at most one) core other than $p$ in the current superstep, i.e., assigning $v$ to $p$ would imply that these $\ell$ vertices can be computed in the next superstep at the earliest. This strategy aims to increase the number of vertices we can compute before having to insert a synchronization barrier. Besides this, to the score of each vertex, we also add the length of the longest outgoing path normalized to a small range $[0,20]$ as a default initialization value, in order to prioritize progress on the critical paths in general. Note that in this scheduler, whenever a vertex is assigned to a core, the priority values on nearby vertices need to be updated to correctly reflect the new number of vertices that would be locked out by each potential assignment. It is thus of dynamic nature.
Our algorithms use dynamically updated priority queues to ensure that the scoring only imposes a minimal overhead on the scheduling.
Besides this, our schedulers are also extended with a fill-up strategy: whenever the algorithm decides that a synchronization barrier is required at time $t$, we specifically check for each core $p$ whether some lower-weight vertices could still fit into the current computation phase on $p$ without delaying the synchronization barrier, and compute as many of these vertices as we can in order of their priority. This is especially useful for DAGs obtained after coarsening, where the compute weights of vertices can differ significantly.
Finally, our schedulers also use a much more sophisticated rule to decide when to close the current superstep. The fraction $\alpha$ of cores that need to be idle to insert a barrier is changed to a parameter $\alpha$, chosen between $0.2$ and $0.4$ for our algorithms. Moreover, our schedulers ensure that even when this threshold is met, our algorithm only inserts a barrier when the current part of the DAG indeed allows for enough parallelization in order to provide work for idle cores.
Altogether, the run time of both of our schedulers is just above linear in the size of the input graph, i.e., in the number of non-zero entries in the matrix. In particular, the worst-case theoretical time complexity is $O(|V|\log|V|+|E|)$ for $p$-ivotal path, and $O(k|V|\log|V|\cdot\sum_{v\in V}{\rm deg}(v)^{2}+|E|)$ for Locking, where $k$ is the number of cores. In practice, our experiments show that the run time of the two schedulers is similar in most cases, and also comparable to that of the HDagg baseline, cf. Section 7.7. In terms of space complexity, $p$-ivotal path requires $O(|V|+|E|)$, whilst Locking requires $O(k|V|+|E|)$ memory.
We provide a detailed pseudocode for our schedulers in Algorithm 3.1, and note:
  • The $\text{ready}_{\text{all}}$ and $\text{ready}_{p}$ structures defined in Lines 3.1, 3.1 of Algorithm 3.1 are priority queues that are always sorted according to the priority values ${\rm prio}[v]$ and ${\rm prio}[p,v]$ of the contained vertices $v$, respectively. It is due to the insertions/deletions from these priority queues that the run time of the schedulers is not strictly linear.
  • In Line 3.1 of Algorithm 3.1, the last condition is specifically that
    $|{\rm ready}_{\rm pool}|\geqslant\min\left(1.2\cdot{\rm busy},\,{\rm busy}+\tfrac{1}{2}\cdot{\rm idle}\right),$ (3.2)
    with ‘${\rm idle}$’ and ‘${\rm busy}$’ denoting the current number of idle and busy cores, respectively. The condition guarantees that inserting a synchronization barrier indeed allows us to employ more cores. Intuitively, the first and second term in the minimum ensure that the increase in parallelism is significant in the cases when ‘${\rm busy}$’ is a small and large value, respectively.
  • Once we decide to insert a barrier (the endStep variable is set), the finishing time $t_{\rm end}$ of the current superstep is finalized as the latest finishing time of vertices that are currently being computed, i.e., the latest entry in finishTimes. Then, until the superstep ends, the fill-up strategy is used: we only start computing a vertex $v$ on a core $p$ if this extra computation is still finished by $t_{\rm end}$ at the latest. In Line 3.1, for each core $p$, we again select $v$ as the highest-priority vertex in $\text{ready}_{p}$ or $\text{ready}_{\text{all}}$ that satisfies this condition.
  • In Line 3.1, the $p$-ivotal path scheduler initializes all vertices of the DAG in linear time using Equation (3.1), via a dynamic programming approach that follows a reverse topological ordering.
  • In contrast to this, in Line 3.1, the Locking scheduler also uses dynamic programming to compute the longest outgoing path from each vertex, then normalizes this to the range $[0,20]$, and uses it as a default priority value. Then, whenever a new vertex is added to the priority queues (in Lines 3.1, 3.1, or 3.1), the rest of the priority function (the number of vertices locked out from the superstep by the assignment) is computed and added to this base priority value.
  • In the Locking scheduler, Line 3.1 considers all children $u$ of $v$ that are not yet locked out of the current superstep. If no parent of $u$ is in the current superstep yet, then for all cores except $p$, the (ready) parents of $u$ get an extra unit of penalty for locking $u$ out from the superstep. If $u$ already has parents on (only one) other core than $p$ in the current superstep, then it is our current assignment that locks $u$ out of the superstep, so all the (ready) parents of $u$ lose a unit of penalty.
@@ 4. Acyclicity-preserving graph coarsening
Previous works discuss several ways to partition a DAG into clusters such that this coarsened graph remains acyclic, although often without a formal proof of this property. In a further generalization of earlier methods from Cong et al. [CLB94, §4] and Zarebavani et al. [ZCL+22, §IV.B], we now introduce the concept of cascades and we prove that coarsening a DAG along such cascades is still guaranteed to preserve acyclicity. This is presented in Section 4.1. In Section 4.2, we describe the graph coarsening algorithm used in our scheduling algorithms.
@@ 4.2. Algorithm
Result: A partition $P$ such that every set $U\in P$ is an in-funnel.
1 Algorithm ${\rm Funnel}(G)$:
@@ 3       ${\rm Visited}[v]\leftarrow\textnormal{{false}},\,\forall v\in V$
6             $U\leftarrow\{v\}$
             $(U,\_)\leftarrow{\rm FunnelDFS}(G,U,v,\emptyset)$
              // see subroutine, Line 4.1
@@ 8                   ${\rm Visited}[u]\leftarrow$ true
11
12      return Partition
13
// Depth-first search invoked by the main algorithm
Data: A directed acyclic graph $G=(V,E)$, a non-empty subset $U\subseteq V$ of vertices, a vertex $u\in U$, and a hash map ${\rm IncludedChildren}$.
Result: A non-empty subset $W\subseteq V$ containing $U$ and a hash map ${\rm IncludedChildren}$.
14 Subroutine ${\rm FunnelDFS}(G,U,u,{\rm IncludedChildren})$ :
15       for $v\in{\rm Parents}(u)$  do
16             if ${\rm IncludedChildren.find}(v)$ then
17                  ${\rm IncludedChildren}[v]\leftarrow{\rm IncludedChildren}[v]+1$
18             else
19                  ${\rm IncludedChildren}[v]\leftarrow 1$
20            if ${\rm IncludedChildren}[v]={\rm OutDegree}(v)$  then
21                   $U.{\rm insert}(v)$
22                   $(U,{\rm IncludedChildren})\leftarrow{\rm FunnelDFS}(G,U,v,{\rm IncludedChildren})$
23
24
25      return $(U,{\rm IncludedChildren})$
26
27
@@ 22                   $(U,{\rm IncludedChildren})\leftarrow{\rm FunnelDFS}(G,U,v,{\rm IncludedChildren})$
In practice, before applying this graph coarsening, we remove some transitive edges from $G$ as this increases the likelihood of finding larger components. A complete transitive reduction is slow, though there are faster approximate transitive reductions, such as the ‘remove all long edges in triangles’-algorithm [PSSD14, §2.3] with a time complexity of $O(\sum_{v\in V}{\rm deg}(v)^{2})$. This algorithm may be terminated early if a faster runtime is desired. In our implementation, we run the full algorithm.
In our implementation, we also add a size/weight constraint on each component of the partition to Algorithm 4.1 as otherwise a graph with only one sink vertex would be coarsened into a graph with only one vertex.
Finally, we remark that we have observed experimentally that coarsening along in-funnels works better than out-funnels. Whether this has something to do with the graphs in the data set or there is an underlying conceptual reason for this is left as an open question.
@@ 5. Reordering for locality
In order to address this, we apply a separate reordering step to ensure that vertices which are computed together are also preferably stored together. The main idea of this approach has already been considered before, cf. [RG92], but has not found its way into modern baselines. In particular, we consider a reordering (relabeling) the vertices of the input DAG based on the partitioning we developed, where we iterate through the supersteps in order, and within each superstep, we iterate through the cores in order. That is, we first start with the vertices $v$ with $\pi(v)=1$, $\sigma(v)=1$, then the vertices $v$ with $\pi(v)=2$, $\sigma(v)=1$, and so on, up to the vertices $v$ with $\pi(v)=k$, $\sigma(v)=1$, followed by vertices $v$ with $\pi(v)=1$, $\sigma(v)=2$, etc. Within a given core-superstep combination, we go through the vertices in the original order (which gives a topological ordering of the induced subDAG). We then symmetrically permute the input matrix and permute the right-hand-side vector of the SpTRSV problem accordingly. Note that since the permutation provides a valid topological ordering of the vertices of the DAG, the resulting matrix is still lower triangular, resulting in an equivalent (symmetrically permuted) formulation of the SpTRSV problem.
@@ 6. Experimental setup
In this section, we present the experimental setup for the evaluation of our scheduling algorithms. Our implementations are available in the OneStopParallel repository [BLPS24] on Github.
@@ 6.1. Methodology
For the evaluation, we used a standard SpTRSV implementation which iterates through the rows of the matrix which was stored in compressed sparse row (CSR) format [TW67]. The algorithm was parallelized using the OpenMP library with the flags OMP_PROC_BIND=close and OMP_PLACES=cores.
We measured one hundred times the time it takes for a single SpTRSV execution using the chrono high-resolution clock. The measurements were taken whilst the system was ‘hot’, meaning an untimed execution precedes the timed executions. Between each SpTRSV execution, the right-hand-side vector $b$ was reset to all ones. The experiments were repeated for each scheduling algorithm, matrix in the data set, and CPU architecture type. The latter two are described in more detail in Section 6.2 and Section 6.3, respectively. If the interquartile range of the measurements corresponding to a scheduling method was too large, we rejected and re-ran all experiments on the same matrix and processor configuration.
@@
All scheduling algorithms are implemented in C++ and were compiled with GCC using the optimization flag -O3.
@@ 6.2. Data sets
For the experiments, we used matrices from several data sets. The main data set is a sample from the SuiteSparse Matrix Collection [DH11], which constitutes a diverse set of matrices from a wide range of applications and was used in previous studies [ZCL+22]. This data set is complemented with two randomly generated ones: uniformly random, i.e., Erdős–Rényi matrices [ER59], and random with a bias towards the diagonal. The former are easier to parallelize as they have few (and thus large) wavefronts [HKSL14] and the latter are designed to be harder to parallelize.
A useful general metric to understand the parallelizability of an SpTRSV execution is the average wavefront size, which can be calculated from the DAG representation by dividing the number of vertices by the length of the longest path. This metric is also indicated for each matrix in the overview of the data sets in the appendices.
@@ 6.2.1. SuiteSparse
We furthermore removed matrices from the data set which had the same sparsity pattern. An overview over some statistics of the matrices may be found in Table LABEL:table:florida-graph-table.
Remark 6.1.
Zarebavani et al. [ZCL+22] also use the real symmetric positive definite matrices from the SuiteSparse Matrix Collection for their data set. Their data set, however, differs from ours insofar as they first symmetrically permute the sparse symmetric matrix using a fill-reducing method of METIS [KK98] and only then take the lower triangular part. In general, this results in non-equivalent SpTRSV problems. As a by-product of this extra step, the matrices in their data set have significantly higher (14×14$\times$14 × geometric-mean) average wavefront size than the lower triangular part of the original matrices, compare Table LABEL:table:hdagg-metis-graph-table with Table LABEL:table:florida-graph-table. As a result, the performance characteristics of HDagg and SpMP reported in their paper differ significantly from our experiments.
We mention that in some contexts, such as with the incomplete Cholesky preconditioned conjugate gradient method for sparse symmetric solve, one is allowed to first symmetrically permute to increase the average wavefront size.

### v2

@@ 3. The GrowLocal scheduler
Rule I: Vertices are prioritized according to
  1.               (i)
    core exclusivity, and then
  2.               (ii)
    smallest ID.
1
2while not all vertices are assigned yet do
3       $\alpha\leftarrow 20$
4       while true  do
             // I. Assign new vertices to each core
5             assign up to $\alpha$ vertices to core $1$ with Rule I
6             $\Omega_{1}\leftarrow$ total newly assigned weight to core $1$
7             for core $p=2,\dots,k$ in order do
8                   $\Omega_{p}\leftarrow 0$
9                   while $\Omega_{p}\not\approx\Omega_{1}$ and can assign to core $p$  do
10                         assign vertex $v$ to core $p$ with Rule I
11                         $\Omega_{p}\leftarrow\Omega_{p}+\omega(v)$
12
13
14
            // II. Check for sufficient parallelism
@@ 11                         $\Omega_{p}\leftarrow\Omega_{p}+\omega(v)$
16            $\beta\leftarrow\frac{\sum_{p}\Omega_{p}}{\max_{p}\Omega_{p}+L}$
17
18            if the parallelization score $\beta$ is high enough then
19                   consider current assignment as worthy
20                   undo new assignments up to the last barrier
21                   $\alpha\leftarrow 1.5\times\alpha$
22
23            else
24                   finalize last worthy assignment as next superstep
25                   break inner loop
@@ 16            $\beta\leftarrow\frac{\sum_{p}\Omega_{p}}{\max_{p}\Omega_{p}+L}$
28
Algorithm 3.1 Skeleton of GrowLocal scheduler
Under reasonably mild assumptions, one can also show that the running time of the algorithm is almost linear.
Theorem 3.1.
Assume that both the out-degrees and compute weights of the DAG are on the same order of magnitude. Then, the time complexity of the GrowLocal algorithm is $O(|E|\cdot\log|V|)$.
The precise formulation of the theorem and the proof are deferred to Section B of the supplementary material. On a high level, since the size of the iterations follows a geometric series, one can show the total number of speculative vertex assignments in a superstep is still only a linear factor more than the size of the finalized superstep. However, a rigorous proof of the theorem is much more technical due to the fact that our parallelization score also depends on the weights of the vertices and the parameter $L$. In the supplementary material, we also provide a brief experimental analysis, which confirms linear complexity.
One can also easily observe that the space requirement of the algorithm is simply $O(|E|)$.
3.1. Block parallel scheduling
In order to reduce the overhead from scheduling even further, one can parallelize the scheduling process. Instead of directly parallelizing the algorithm, we split the scheduling problem into independent parts. We achieve this by subdividing the lower triangular matrix into smaller lower triangular matrix blocks along the diagonal as in Figure 3.1.
(a) Decomposition of $9\times 9$ matrix
DAG $1$DAG $2$DAG $3$
(b) Sub-DAGs
Figure 3.1. Subdivision of a $9\times 9$ lower triangular matrix into three $3\times 3$ lower triangular matrix blocks (a) and the three corresponding sub-DAGs with inter-DAG dependency (b).
On each of these sub-problems, we can generate schedules in parallel. When combining these schedules, we just have to ensure that the individual schedules are combined one after the other. Equivalently, we can add to the superstep assignment of each vertex in each block the total number of supersteps in earlier block schedules.
We remark that for the weight of the vertices (in the DAG representation), we still use the number of non-zeros in the full matrix. This is in line with our SpTRSV kernel implementation.
@@ 4. Acyclicity-preserving graph coarsening
Previous works discuss several ways to partition a DAG into clusters such that this coarsened graph remains acyclic, although often without a formal proof of this property. In a further generalization of earlier methods from Cong et al. [CLB94, §4] and Zarebavani et al. [ZCL+22, §IV.B], we now introduce the concept of cascades and prove that coarsening a DAG along such cascades is still guaranteed to preserve acyclicity. This is presented in Section 4.1. In Section 4.2, we describe the graph coarsening algorithm used in our scheduling algorithms.
@@ 4.2. Algorithm
Result: A partition $P$ such that every $U\in P$ is an in-funnel.
@@ 2 ${\rm Visited}[v]\leftarrow\textnormal{{false}},\,\forall v\in V$
5       $U\leftarrow\emptyset$
6       ${\rm ChildrenCount}[u]\leftarrow 0,\,\forall u\in V$
7       ${\rm PrioQueue}.{\rm insert}(v)$
8       while not ${\rm PrioQueue}.{\rm empty}()$ do
9             $w\leftarrow{\rm PrioQueue}.{\rm pop}()$
10             $U.{\rm insert}(w)$
11             for $u\in{\rm Parents}(w)$  do
12                   ${\rm ChildrenCount}[u]\leftarrow{\rm ChildrenCount}[u]+1$
13                   if ${\rm ChildrenCount}[u]={\rm OutDegree}(u)$  then
14                         ${\rm PrioQueue}.{\rm insert}(u)$
15
16
17
@@ 19             ${\rm Visited}[u]\leftarrow$ true
22
23return Partition
@@
In practice, before applying this graph coarsening, we remove some transitive edges from $G$ as this increases the likelihood of finding larger components. A complete transitive reduction is slow, though there are faster approximate transitive reductions, such as the ‘remove all long edges in triangles’-algorithm [PSSD14, §2.3] with a time complexity of $O(\sum_{v\in V}{\rm deg}(v)^{2})$. This algorithm may be terminated early if a faster runtime is desired. In our implementation, we run the full (approximate) algorithm.
In our implementation, we also add a size/weight constraint on each component of the partition to Algorithm 4.1 as otherwise a graph with only one sink vertex would be coarsened into a graph with only one vertex. Altogether, our coarsening algorithm Funnel is a more general [^2] Every in-tree is an in-funnel. and robust version of the coarsening algorithm in HDagg [ZCL+22].
@@ 5. Reordering for locality
In order to address this, we apply a separate reordering step to ensure that vertices which are computed together are also stored together. The main idea of this approach has already been considered before, cf. [RG92], but has not found its way into modern baselines. In particular, we consider a reordering (relabeling) the vertices of the input DAG based on the partitioning we developed, where we iterate through the supersteps in order, and within each superstep, we iterate through the cores in order. That is, we first start with the vertices $v$ with $\pi(v)=1$, $\sigma(v)=1$, then the vertices $v$ with $\pi(v)=2$, $\sigma(v)=1$, and so on, up to the vertices $v$ with $\pi(v)=k$, $\sigma(v)=1$, followed by vertices $v$ with $\pi(v)=1$, $\sigma(v)=2$, etc. Within a given core-superstep combination, we go through the vertices in the original order (which gives a topological ordering of the induced sub-DAG). We then symmetrically permute the input matrix and permute the right-hand-side vector of the SpTRSV problem accordingly. Note that since the permutation provides a valid topological ordering of the vertices of the DAG, the resulting matrix is still lower triangular, resulting in an equivalent (symmetrically permuted) formulation of the SpTRSV problem.
@@ 6. Experimental setup
In this section, we present the experimental setup for the evaluation of our scheduling algorithm. Our implementations are available in the OneStopParallel repository [BLM+24] on Github.
@@ 6.1. Methodology
For the evaluation, we used a standard SpTRSV implementation which iterates through the rows of the matrix which was stored in compressed sparse row (CSR) format [TW67]. The algorithm was parallelized using the OpenMP library with the flags OMP_PROC_BIND and OMP_PLACES set to close and cores, respectively.
We measured one hundred times the time it takes for a single SpTRSV execution using the chrono high-resolution clock. The measurements were taken whilst the system was ‘hot’, meaning two untimed executions precede the timed executions. Between each SpTRSV execution, the right-hand-side vector $b$ was reset to all ones. The experiments were repeated for each scheduling algorithm, data set, and CPU architecture type. The latter two are described in more detail in Section 6.2 and Section 6.3, respectively. If the interquartile range of the measurements corresponding to a scheduling method was too large, we rejected and re-ran all experiments on the same matrix and processor configuration.
@@
All scheduling algorithms are implemented in C++ and were compiled with GCC (11.4.0 or 11.5.0) using the optimization flag -O3.
@@ 6.2. Data sets
For the experiments, we used matrices from several data sets. The main data set is a sample from the SuiteSparse Matrix Collection [DH11], which constitutes a diverse set of matrices from a wide range of applications and was used in previous studies [ZCL+22]. We also consider two modified versions of this data set that are also relevant in their own right. Finally, these data sets are complemented with two randomly generated ones: uniformly random, i.e., Erdős–Rényi matrices [ER59], and random with a bias towards the diagonal. The former are easier to parallelize as they have few (and thus large) wavefronts [HKSL14] and the latter are specifically designed to be harder to parallelize, though they admit good locality.
A useful general metric to understand the parallelizability of an SpTRSV execution is the average wavefront size, which can be calculated from the DAG representation by dividing the number of vertices by the length of the longest path. This metric is indicated for each matrix in the overview of the data sets in the supplement, see Section A.
@@ 6.2.1. SuiteSparse
We furthermore removed matrices from the data set which had the same sparsity pattern. An overview over some statistics of the matrices may be found in Table LABEL:table:florida-graph-table of the supplement.
6.2.2. SuiteSparse METIS (METIS)
Zarebavani et al. [ZCL+22] also use the real symmetric positive definite matrices from the SuiteSparse Matrix Collection as their data set. However, they use a modified version of this data set, which we also reproduce here. In their experiments, the matrices are first symmetrically permuted using a fill-reducing method of METIS [KK98] and only then the lower triangular part is taken. In general, this results in non-equivalent SpTRSV problems. The sparsity pattern of the matrices in this data set are representative of SpTRSV workloads in a Gauß–Seidel or a zero-fill-in incomplete Cholesky preconditioned conjugate gradient method for sparse symmetric solve. An overview over some statistics of the matrices may be found in Table LABEL:table:hdagg-metis-graph-table of the supplement.
6.2.3. SuiteSparse Eigen incomplete Cholesky (iChol)
This data set consists of lower triangular matrices obtained after an incomplete Cholesky decomposition. The initial set of matrices are the same symmetric matrices used in the SuiteSparse data set [^4] The matrix ‘bundle_adj’ segmentation-faults during the process and is thus excluded from the data set.. The incomplete Cholesky decomposition was performed using the ‘IncompleteCholesky’ method of Eigen [GJ+10] using the built-in fill-reducing method ‘AMDOrdering’. An overview over some statistics of the matrices may be found in Table LABEL:table:cholesky-graph-table of the supplement.